40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Feb 20 16:20:33 2021, gautam, Update, General, Housekeeping + PRMI char IPPOSdeat.pdfTTdrift.pdfPOXnPOY.pdfPRMI.pdf
    Reply  Sun Feb 21 20:51:21 2021, rana, Update, General, Housekeeping + PRMI char 
       Reply  Tue Feb 23 00:10:05 2021, gautam, Update, General, Demod char part 1 noPreamp.pdfwithPreamp.pdf
          Reply  Wed Feb 24 11:53:24 2021, gautam, Update, General, Demod char part 2 noises_noPreamp.pdfnoises_withPreamp.pdf
             Reply  Wed Feb 24 12:11:08 2021, gautam, Update, General, Demod char part 3 noPreamp.pdfwithPreamp.pdfnoPreamp_whitened.pdfwithPreamp_whitened.pdf
                Reply  Thu Feb 25 16:50:53 2021, gautam, Update, General, PRMI sensing matrix PRMI1f_noArmssensMat.pdf
                   Reply  Thu Feb 25 20:37:49 2021, gautam, Update, General, Setting modulation frequency and checking IMC offset modFreq.pdfIMC_offset.pdf
Message ID: 15844     Entry time: Thu Feb 25 16:50:53 2021     In reply to: 15840     Reply to this: 15845
Author: gautam 
Type: Update 
Category: General 
Subject: PRMI sensing matrix 

After all the work at the LSC rack over the last couple of days, I re-locked the PRMI (ETMs misaligned), and measured the sensing matrix once again. The PRMI was locked using 1f error signals, with AS55_Q as the MICH sensor and REFL11_I as the PRCL sensor. As shown in Attachment #1, the situation has not changed, there is still no separation between the DoFs in the REFL signals. I will measure the MC lock point offset using the error point dither technique today to see if there is something there.

Attachment 1: PRMI1f_noArmssensMat.pdf  38 kB  | Hide | Hide all
PRMI1f_noArmssensMat.pdf
ELOG V3.1.3-